OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 4) and a directrix at y = -4. Answer options: y = 1 /16x2 y^2 = 16x y^2 = 4x y = 1/ 4x2

OpenStudy (mathmate):

If the focus is at (0,4), and the directrix is at y=-4, the parabola opens upwads, so the standard equation is \(\large y-k=\frac{(x-h)^2}{4p}\) where (h,k) is the position of the vertex. It so happens that the vertex is half-way between the focus and the directrix, and the half distance is p. 

OpenStudy (mathmate):

So your first step is find (h,k), the point half-way between focus and the directrix. then substitute in the standard formula and simplify.

OpenStudy (mathmate):

It turns out that (h,k)=(0,0) because x=0, y=(4+-4)/2=0. That should make your life a little easier.